OpenStudy (anonymous):

Write in order from the least to greatest 3^2, -l7l, -(-10) , - l1l ?

OpenStudy (anonymous):

Do not do the equation.

OpenStudy (xishem):

3^2 = 9 -171 = -171 -(-10) = 10 -111 = -111 Are these the numbers you wrote? It looks like you used l's (L).

OpenStudy (anonymous):

-171, -111, 3^2, 10

OpenStudy (anonymous):

Yes I couldn't find the spacing things, so I used l's.

OpenStudy (xishem):

Why couldn't you just use 1's (ones). Anyway, Sebivid is correct.